Geometry. The length of a rectangle is 1 cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions (the length and the width)of the rectangle?
Help would be much appreciated! Thank you.

Answer by scott8148(6628)

let x=width, so x+1=length ... the width, length and diagonal form a right triangle

by Pythagoras, x^2+(x+1)^2=4^2 ... x^2+x^2+2x+1=16 ... 2x^2+2x-15=0